Hotspur Corp. has 50 million shares that are currently trading at $4 per share, and $200 million worth of debt. The debt is risk free and has interest rate of 5% and expected return of stock is 11% for this company.

The price of the stock falls to $3 per share but there is no change to the value of risk-free debt or risk (asset beta) of their assets. We are in Modligliani-Miller world (no taxes).

1- What happens to their equity cost of capital (rE)

2- Using the new rE, would you suggest a change in capital structure to the CFO? Would you suggest higher or lower leverage and why?
Answer #1

WACC = rd * D/ V + re * E/V

= 5% * ( 200/ 200+( 50*4)) + 11%* ( 50*4/ 200+( 50*4))

= 5% * 0.5 + 11% * 0.5 = 8%

When stock price changed to $3 , value of equity changes and thus capital struxture changes . To adjust that cost of equity will also change keeping wacc constant as per modigliani miller theory 2 without taxes.

WACC = rd * D/ V + re * E/V

8% = 5% * ( 200/ 200+( 50*3)) + re * ( 50*3/ 200+( 50*3))

8% = 5%* 2/3.5 + re * 1.5/3.5

re = 12%

To decrease re , leverage needs to be decreased